Grade 12Physical Chemistry

At 25°C and 760 mm of Hg pressure a gas occupies 600 mL volume. What will be its pressure at a height where temperature is 10°C and volume of the gas is 640 mL ?

sudhanshu
12 Years agoGrade 12

4 Answers

Gaurav
11 Years ago
Hello Student
P1 = 760 mm Hg
V1 = 600 Ml
T1 = 298 K
V2 = 640 Ml
T2 = 283 K
Using
P1V1= P2V2
We get
P2 = 676.6 mm Hg
Sushrut
7 Years ago
P1=760 mm Hg
V1=600 mL
T1=298k
V2=640mL
T2=283k
P2= ?
       Now,
        We know that;
        P1 V1 = P2 V2 is same as P1 V1/T1 = P2 V2/T2
      =760×600/298=640×P2/283
      =1530.20/2.261
      =676mm Hg
 
         P1V1=P2V2 is same as P1V1/T1 =P2V2
Hahahaha
7 Years ago
Hahsaayayaayadayadyayadydayadydsayadydsaayadydsaaayadydsaaaayadydsaaaaayadydsaaaaaayadydsaaaaaaayadydsaaaaaaaayadydsaaaaaaaà
 
Rishi Sharma
5 Years ago
Dear Student,
Please find below the solution to your problem.

P1 = 760 mm Hg
V1 = 600 Ml
T1 = 298 K
V2 = 640 Ml
T2 = 283 K
Using
P1V1= P2V2
We get
P2 = 676.6 mm Hg

Thanks and Regards